OpenStudy (anonymous):

Find the value of log√7^1.

OpenStudy (anonymous):

\[\sqrt{7}^1\]

OpenStudy (jhannybean):

\[\log_{10} (\sqrt{7})^{1}\]?

OpenStudy (anonymous):

I'm not completely sure. Thats all the directions say on my homework.

OpenStudy (jhannybean):

Is the base 10? Because if a value is not given the base of logs is usually 10.

OpenStudy (anonymous):

It does not say.

OpenStudy (camper4834):

wait wait

OpenStudy (camper4834):

\[\log_{\sqrt{7}}1 \]

OpenStudy (camper4834):

does it look like that?

OpenStudy (jhannybean):

Well, let's see... \(\sqrt{7} = 7^{1/2}\)

OpenStudy (anonymous):

Yes!

OpenStudy (camper4834):

OOOOKAY its base sqrt(7)

OpenStudy (anonymous):

i bet $2 that that is the question

OpenStudy (anonymous):

it looks exactly like camper4834 said.

OpenStudy (jhannybean):

So this could be written as \[\log(7^{1/2})\]

OpenStudy (anonymous):

makes no difference at all what the base is \[\log_b(1)=0\] since \[b^0=1\] always

OpenStudy (camper4834):

\[\sqrt{7}^x = 1\]

OpenStudy (jhannybean):

Oh. in that case \[\log_{\sqrt{7}} 1 = x\] Now you're asking yourself \(\sqrt{7}^x =1\)

OpenStudy (jhannybean):

so.. \(\sqrt{7} = 7^{1/2 \cdot x}\)

OpenStudy (anonymous):

So 7^1/2x is the answer?

OpenStudy (jhannybean):

Nope.

OpenStudy (anonymous):

Okay can you help me get to the next step?

OpenStudy (anonymous):

at the risk of repeating myself \[\huge \log_b(1)=0\] always

OpenStudy (anonymous):

because irrespective of the base, \[b^0=1\]

OpenStudy (camper4834):

Bcolllins listen to satellite

OpenStudy (camper4834):

what he is saying is that the answer is ZERO

OpenStudy (anonymous):

that isn't the answer.

OpenStudy (jhannybean):

Oh, I guess I made it too complicated. <.<

OpenStudy (anonymous):

0 is not the answer.

OpenStudy (anonymous):

Nevermind. Thank you so much satellite!

OpenStudy (anonymous):

\[\log_{\sqrt{7}}(1)=0\]for sure, maybe i did not understand the question

OpenStudy (anonymous):

Yes, it is the answer, I had typed a space and then a 0 and it was counting it wrong. Thank you so much!!

OpenStudy (jhannybean):

Ah, then I was just confused. I was thinking something along the lines of....

OpenStudy (anonymous):

Would either of ya'll know how to do........ evaluate ln e^6 without a calculator?

OpenStudy (jhannybean):

\[\log_{\sqrt{7}} 1=x\]\[\sqrt{7}^x = 1\]\[7^{(1/2)x} = 1\]\[\log_{7} 7^{(1/2)x} = \log_{7} 1\] Which would just get way too messy and complicated.

OpenStudy (jhannybean):

That's what I thought it was at first, I had just read the question wrong. sorry :')